# Variable Selection in Convex Piecewise Linear Regression

### v1

@@ 3.2 Theoretical Analysis of Initialization
Ghosh et al. [2021] showed that their spectral initialization provides an $\epsilon$-accurate subspace estimation with $\widetilde{\mathcal{O}}(\epsilon^{-2}d)$ samples (when the ground-truth and hence the model order $k$ are fixed). In the sparse case of max-affine regression, Theorem 3.1 reduces the dependence of the sample complexity on $d$ from linear to logarithmic. Note that the spectral gap $\delta_{\mathrm{gap}}$ in Theorem 3.1 is also independent of $d$ similar to [Ghosh et al., 2021, Theorem 2].
@@
We will use the following known results about $\bm{M}$: i) The column space of $\bm{M}$ coincides with the $k$-dimensional subspace spanned by the ground-truth weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ [Ghosh et al., 2021, Lemma 3]; ii) There exists $\delta_{\mathrm{gap}}>0$, independent of $d$, such that the smallest nonzero eigenvalue of $\bm{M}$ is bounded from below by $\delta_{\mathrm{gap}}$ [Ghosh et al., 2021, Lemma 7]. Then [Vu et al., 2013, Theorem 3.1] provides a perturbation bound given by
$\|\widehat{\bm{P}}-\bm{P}\|_{\mathrm{F}}\leq\frac{4s}{\delta_{\mathrm{gap}}}\cdot{\left\lVert\bm{\widehat{M}}-\bm{M}\right\rVert}_{\infty}.$ (26)
@@
We further proceed with the remainder of the proof under the event that (28) holds. The first assertion $\mathcal{S}=\mathcal{S}^{\star}$ follows from [Vu et al., 2013, Theorem 3.2] if
@@
Once the support $\mathcal{S}^{\star}$ is exactly recovered by Algorithm 2, the estimation accuracy of Algorithm 3 is provided by [Ghosh et al., 2021, Theorem 3] with $d$ substituted by $s$. We provide the statement of this result in our notation for completeness.
Theorem 3.3 (A paraphrase of [Ghosh et al., 2021, Theorem 3]).
Instate the assumptions in Theorem 3.1. Let $R_{\max}:=\max_{j\in[k]}{\left\lVert\bm{\theta}^{\star}_{j}\right\rVert}$. Then it holds with probability at least $1-n^{-11}$ that the initial parameter estimate $\{\bm{\theta}^{0}_{j}\}_{j=1}^{k}$ by Algorithm 3 with $\mathcal{S}=\mathcal{S}^{\star}$ satisfies
@@
The estimation error in Theorem 3.3 is stated using the minimum distance between the ground-truth parameters $\{\bm{\theta}^{\star}_{j}\}_{j=1}^{k}$ and all permutations of the estimates $\{\bm{\theta}^{0}_{j}\}_{j=1}^{k}$. On the other hand, Ghosh et al. [2021, Theorem 3] stated the estimation error as the minimum distance up to both permutation and scaling ambiguities. However, careful examination of their proof shows that their error bound applies to a particular scaling with the minimizer $c^{\sharp}$ in Algorithm 3. Therefore, [Ghosh et al., 2021, Theorem 3] also implies (31).
@@
Instate the assumptions of Theorems 3.1 and 3.3. Let $\epsilon\in(0,1)$. Then it holds with probability at least $1-n^{-11}$ that applying Algorithms 2 and 3 in succession yields an initial estimate satisfying
@@
Theorem 3.5 reduces the sample complexity of the spectral initialization for max-affine regression by Ghosh et al. [2021] when the weight vectors satisfy the joint $s$-sparsity structure. Specifically, the linear dependence on the ambient dimension $d$ drops to polynomial dependence on $s$, significantly reducing the sample complexity when $s\ll d$. The order of the polynomial in $s$ depends on the geometry of the model parameters $\{\bm{\theta}^{\star}_{j}\}_{j=1}^{k}$. For example, if $\bm{1}\in\mathbb{R}^{s}$ is in the span of $\{\bm{\theta}^{\star}_{j}\}_{j=1}^{k}$, then it can be shown that $\min_{j\in[s]}[\bm{P}]_{jj}\geq 1/s$, whence the sample complexity in Theorem 3.5 becomes $\mathcal{O}(s^{4})$.
@@ 4.1 Phase Transitions of Sp-GD
This subsection presents numerical results of the Sp-GD algorithm that corroborate the theoretical guarantees presented in Section 2.2. In the simulation, we initialized Sp-GD in two steps. First, we estimate the parameter subspace by Algorithm 2. Second, we apply a practical alternative to Algorithm 3 which randomly samples from the estimated subspace and chooses the best one that produces the smallest fit error after $10$ iterations of Sp-GD. We adopt this heuristic instead of Algorithm 3 for the following reasons. The first step of Algorithm 3 creates an $r$-covering of the unit $\ell_{2}$-ball $B^{k+1}$, where its cardinality scales as $\mathcal{O}(r^{-k})$ [Vershynin, 2018, Corollary 4.2.12]. Therefore, the exhaustive search in the second step over all elements in the $r$-covering becomes impractical as the parameter $r$ decreases. However, the accuracy of Algorithm 3 crucially depends on $r$. The estimation performance is evaluated via the median of the relative error between the true model coefficients $\bm{\theta}^{\star}\triangleq(\bm{\theta_{j}}^{\star})_{j=1}^{k}$ and the estimated coefficients $\widehat{\bm{\theta}}\triangleq(\widehat{\bm{\theta}}_{j})_{j=1}^{k}$ over $50$ Monte Carlo simulations. The relative error is defined via the optimal permutation of the affine model indices as
@@
Figure 1: Median of $\mathrm{err}(\widehat{\bm{\theta}})$ for different ($n$,$d$) pairs using 50 Monte Carlo iterations for $k=3$ and $s=25$ with Gaussian (top) and Uniform (bottom) covariate distributions in the noiseless case.
@@
Figure 2: Median of $\mathrm{err}(\widehat{\bm{\theta}})$ for different ($n$,$s$) pairs using 50 Monte Carlo iterations for $k=3$ and $d=200$ with Gaussian (left) and Uniform (right) covariate distributions. The red curves are fitted with respect to $s\log(d/s)$ at the phase transition boundary for both figures.
Figure 3: Median of $\mathrm{err}(\widehat{\bm{\theta}})$ for different ($n$,$\sigma_{z}^{2}$) pairs using 50 Monte Carlo iterations for $s=50$, $d=200$ and $k=3$ with Gaussian covariates and local initial estimate.
@@ 4.2 Subspace Estimation and Random Search for Initialization
This subsection delves into the detailed empirical analysis of the initialization method employed in the previous section. First, Fig. 4 demonstrates the gain of Algorithm 2 with SPCA over the analogous spectral method with PCA. The regularization parameter $\lambda>0$ required by Algorithm 2 is set by a parameter sweep over a sampling grid. Alternatively, this parameter can be tuned through cross-validation.
@@
Next, we investigate the empirical performance of the repeated random initialization that replaces the exhaustive discrete search in Algorithm 3 by random sampling followed by $10$ iterations. In particular, we observe the estimation accuracy as the function in the number of random initializations $M$. Fig. 5 compares the estimation error by the repeated random initialization when the parameter subspace is estimated by the PCA-based spectral initialization [Ghosh et al., 2021] (red) and by Algorithm 2 (blue). For both of the two subspace estimation methods, the initialization error monotonically decreases as the number of trials $M$ increases. However, the initialization performance by Algorithm 2 is more accurate since it provides better subspace estimation performance.
@@ 5 Discussion
We consider variable selection for a class of non-linear regression models given by the maximum of $k$ affine models $\bm{x}\mapsto\max_{j\in[k]}\langle\bm{a}_{j}^{\star},\bm{x}\rangle+b_{j}^{\star}$ for $j=1,\dots,k$ where $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ and $\{b_{j}^{\star}\}_{j=1}^{k}$ denote the ground-truth weight vectors and intercepts. The weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ satisfy the joint $s$-sparse structure as we assume that only $s$ out of $d$ covariates contributes to explaining the response variable. This paper proposes a variant of projected gradient algorithm called Sp-GD to estimate the sparse model parameters. We provide non-asymptotic local convergence guarantees for Sp-GD under independent sub-Gaussian noise when the covariates follow a sub-Gaussian distribution satisfying the anti-concentration property. Under these assumptions, when the ground-truth model order and parameters are fixed, a suitably initialized Sp-GD converges linearly to an $\epsilon$-accurate parameter estimate given $\mathcal{O}(\max(\epsilon^{-2}\sigma_{z}^{2},1)s\log(d/s))$. In particular when the observations are noise-free ($\sigma_{z}^{2}=0$), Sp-GD guarantees exact parameter recovery.
@@
Since minimizing the squared loss of sparse max-affine models is non-convex, starting Sp-GD within the basin of attraction is crucial for its convergence to the desired estimate. For this purpose, we propose a modification of the spectral method by Ghosh et al. [2021] which estimates the span of the max-affine weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ so that the subspace estimation utilizes the jointly sparse structure in the weight vectors via sparse principal component analysis. Combined with the $r$-covering search over the estimated subspace, the initialization scheme provides an $\epsilon$-accurate estimate when $r=\mathcal{O}(\epsilon)$ given $\mathcal{O}(\epsilon^{-2}\max(\sigma_{z}^{4},\sigma_{z}^{2},1)s^{2}\log^{4}d)$ observations when the ground-truth model parameters are fixed, and the covariates and noise follow Gaussian distributions. The dominating factor of the sample complexity is $s^{2}$, which is significantly smaller than $d$ in the non-sparse case. One noteworthy limitation of the initialization is the $r$-covering search inherited from the previous work on non-sparse max-affine regression [Ghosh et al., 2021]. The cost of constructing the $r$-covering and the exhaustive search over it increases exponentially in the subspace dimension. Therefore, the $r$-covering search is not practical for small separation parameter $r$ which is necessary for accurate estimates. However, this is the only known algorithm with theoretical guarantees for max-affine parameter initialization. Therefore, it would be a fruitful future direction to develop a practical initialization scheme that avoids the $r$-covering search and provides theoretical performance guarantees. Alternatively, it would also be intriguing to explore potential analysis of Sp-GD from random initialization that will extend the known theoretical results on single-index models [Tan and Vershynin, 2019, Chandrasekher et al., 2022].
@@
$\left\|\bm{\theta}^{t+1}-\bm{\theta}^{\star}\right\|^{2}_{2}\leq\tau\left\|\bm{\theta}^{t}-\bm{\theta}^{\star}\right\|^{2}_{2}+\underbrace{C_{1}\sigma^{2}_{z}\left(\frac{sk\log\left(n/s\right)+s\log\left({d}/{s}\right)+\log\left({1}/{\delta}\right)}{n}\right)}_{\rho_{\mathrm{noise}}},$ (38a)
$\bm{\theta}^{t+1}\in\mathcal{N}(\bm{\theta}^{\star},\sqrt{2}\Delta\rho)$ (38b)
for some absolute constant $C_{1}$, hold for all $t\in\mathbb{N}\cup\{0\}$. We prove this statement by induction. Let $t$ be arbitrarily fixed. We assume that (38) holds for all previous iterates.
To show that (38a) also holds for the current iterate, we introduce the following notation. Let $\mathcal{S}^{\star}$ and $\mathcal{S}^{t}$ denote the joint support of $\{\bm{\theta}_{j}^{\star}\}_{j=1}^{k}$ and $\{\bm{\theta}_{j}^{t}\}_{j=1}^{k}$, respectively. The union of $\mathcal{S}^{\star}$ and $\mathcal{S}^{t}$ is denoted by $\varphi^{t}$. The step-size used in Sp-GD stated in Algorithm 1 is stated as
@@
$\left\|\bm{\theta}^{t+1}-\bm{\theta}^{\star}\right\|_{2}$ $=\left\|\left(\bm{I}_{k}\otimes\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right)\left(\bm{\theta}^{t+1}-\bm{\theta}^{\star}\right)\right\|_{2}$
$\leq\left\|\left(\bm{I}_{k}\otimes\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right)\left(\bm{\theta}^{t+1}-\bm{\alpha}^{t+1}\right)\right\|_{2}+\left\|\left(\bm{I}_{k}\otimes\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right)\left(\bm{\alpha}^{t+1}-\bm{\theta}^{\star}\right)\right\|_{2}$
$=\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\theta}^{t+1}_{j}-\bm{\alpha}^{t+1}_{j}\right)\right\|_{2}^{2}\right)^{1/2}+\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\alpha}^{t+1}_{j}-\bm{\theta}^{\star}_{j}\right)\right\|_{2}^{2}\right)^{1/2}$
$\leq 2\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\alpha}^{t+1}_{j}-\bm{\theta}^{\star}_{j}\right)\right\|_{2}^{2}\right)^{1/2}=2\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\theta}^{t}_{j}-\mu_{j}^{t}\nabla_{\bm{\theta}_{j}}\ell(\bm{\theta}^{t})-\bm{\theta}^{\star}_{j}\right)\right\|_{2}^{2}\right)^{1/2},$ (40)

### v2

@@ 3.2 Theoretical Analysis of Initialization
The work in (Ghosh et al., 2021) showed that their spectral initialization provides an $\epsilon$-accurate subspace estimation with $\widetilde{\mathcal{O}}(\epsilon^{-2}d)$ samples (when the ground-truth and hence the model order $k$ are fixed). In the sparse case of max-affine regression, Theorem 3.1 reduces the dependence of the sample complexity on $d$ from linear to logarithmic. Note that the spectral gap $\delta_{\mathrm{gap}}$ in Theorem 3.1 is also independent of $d$ similar to (Ghosh et al., 2021, Theorem 2).
@@
We will use the following known results about $\bm{M}$: i) The column space of $\bm{M}$ coincides with the $k$-dimensional subspace spanned by the ground-truth weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ (Ghosh et al., 2021, Lemma 3); ii) There exists $\delta_{\mathrm{gap}}>0$, independent of $d$, such that the smallest nonzero eigenvalue of $\bm{M}$ is bounded from below by $\delta_{\mathrm{gap}}$ (Ghosh et al., 2021, Lemma 7). Then (Vu et al., 2013, Theorem 3.1) provides a perturbation bound given by
$\|\widehat{\bm{P}}-\bm{P}\|_{\mathrm{F}}\leq\frac{4s}{\delta_{\mathrm{gap}}}{\left\lVert\bm{\widehat{M}}-\bm{M}\right\rVert}_{\infty}.$ (28)
@@
We further proceed with the remainder of the proof under the event that (30) holds. The first assertion $\mathcal{S}=\mathcal{S}^{\star}$ follows from (Vu et al., 2013, Theorem 3.2) if
@@
Once the support $\mathcal{S}^{\star}$ is exactly recovered by Algorithm 2, the estimation accuracy of Algorithm 3 is provided by (Ghosh et al., 2021, Theorem 3) with $d$ substituted by $s$. We provide the statement of this result in our notation for completeness.
Theorem 3.3 (A paraphrase of (Ghosh et al., 2021, Theorem 3)).
Instate the assumptions in Theorem 3.1. Let $R_{\max}\triangleq\max_{j\in[k]}\|\bm{\theta}^{\star}_{j}\|$. Then it holds with probability at least $1-n^{-11}$ that the initial parameter estimate $\{\bm{\theta}^{0}_{j}\}_{j=1}^{k}$ by Algorithm 3 with $\mathcal{S}=\mathcal{S}^{\star}$ satisfies
@@
The estimation error in Theorem 3.3 is stated using the minimum distance between the ground-truth parameters $\{\bm{\theta}^{\star}_{j}\}_{j=1}^{k}$ and all permutations of the estimates $\{\bm{\theta}^{0}_{j}\}_{j=1}^{k}$. On the other hand, (Ghosh et al., 2021, Theorem 3) stated the estimation error as the minimum distance up to both permutation and scaling ambiguities. However, careful examination of their proof shows that their error bound applies to a particular scaling with the minimizer $c^{\sharp}$ provided by Algorithm 3. Therefore, (Ghosh et al., 2021, Theorem 3) also implies (33).
@@
Instate the assumptions of Theorems 3.1 and $3.$ Let $\epsilon\in(0,1)$. Then it holds with probability at least $1-n^{-11}$ that applying Algorithms 2 and 3 in succession yields an initial estimate satisfying
@@
Theorem 3.5 reduces the sample complexity of the spectral initialization for max-affine regression by the method in (Ghosh et al., 2021) when the weight vectors satisfy the joint $s$-sparsity structure. Specifically, the linear dependence on the ambient dimension $d$ drops to polynomial dependence on $s$, significantly reducing the sample complexity when $s\ll d$. The order of the polynomial in $s$ depends on the geometry of the model parameters $\{\bm{\theta}^{\star}_{j}\}_{j=1}^{k}$. For example, if $\bm{1}\in\mathbb{R}^{s}$ is in the span of $\{\bm{\theta}^{\star}_{j}\}_{j=1}^{k}$, then it can be shown that $\min_{j\in[s]}[\bm{P}]_{jj}\geq 1/s$, hence the sample complexity in Theorem 3.5 becomes $\mathcal{O}(s^{4})$.
4 Learning Sparse Generalized Polynomials Via Max-Affine Regression
In this section, we present theoretical guarantees for the dequantization error by Real Maslov Dequantization (RMD) defined in (6), and the non-asymptotic convergence guarantee for Sp-GD under the bounded additive noise model. For convenience, we restate the RMD transformation here as
$y=\mathrm{Re}\{\varsigma\log w\},\quad x_{l}=\varsigma\log u_{l},\quad\forall l\in[d].$ (40)
In what follows, collect the exponents of the generalized polynomial model defined in (4) as $\bm{\alpha}^{\star}_{j}=[\alpha^{\star}_{j,1};\ldots;\alpha^{\star}_{j,d}]$ for every $j\in[k]$. The dequantization error by RMD with temperature parameter $\varsigma>0$ is written as
$z_{\varsigma}:=\max_{j\in[k]}\langle[x_{1};\ldots;x_{d};1],[\bm{\alpha}^{\star}_{j};\log|c^{\star}_{j}|]\rangle-y.$ (41)
The following theorem provides a uniform bound for the dequantization error by RMD.
Theorem 4.1.
Given the generalized sparse polynomial relation $w=g(u_{1},\ldots,u_{d})$ defined in (4), let $y=\mathrm{Re}\{\varsigma\log w\}$ and $\bm{x}=\varsigma\log\bm{u}$ be the transformed variables with $\varsigma>0$. If $\bm{x}$ satisfies Assumption 2.2 and $\varsigma>0$ is sufficiently small, then it holds with probability at least $1-(\gamma\vartheta)^{\zeta}$ that
$|z_{\varsigma}|\leq 2\varsigma(k-1)\exp\left(\frac{-\sqrt{\vartheta}\Delta}{\varsigma}\right),$ (42)
where $\Delta>0$ is the minimum separation parameter defined in (15).
The proof of this theorem is deferred to Appendix E. Theorem 4.1 shows that RMD has a dequantization error decaying exponentially in $\Delta/\varsigma$. This implies that for the dequantization error to be bounded as $|z_{\varsigma}|\leq\epsilon$, then we need to select $1/\varsigma\geq\mathcal{O}(\log\frac{k}{\epsilon})$. Therefore, via RMD, we can learn the real exponents $\alpha^{\star}_{j,l}$’s through sparse max-affine regression. By plugging in the exponent estimates into (4), the coefficients $\{c_{j}^{\star}\}_{j=1}^{k}$ can be easily approximated by linear least squares. We note that the dequantization error bound in Theorem 4.1 has not been shown even in the original Maslov dequantization for the simpler posynomial model by (Maragos et al., 2021).
Building on the non-asymptotic theory in Theorem 2.3, we can show the analysis of estimating the exponents in (4). This requires modifying the proof of Theorem 2.3 to handle the bounded noise model of $z_{\varsigma}$ that is also dependent on the covariate $\bm{x}$.
Corollary 4.2.
Suppose that $\{\bm{u}_{i}\}_{i=1}^{n}$ are independent copies of $\bm{u}$ is distributed such that $\bm{x}=\varsigma\log\bm{u}$ satisfies Assumptions 2.12.2. Let the targets $\{w_{i}\}_{i=1}^{n}$ be generated according to (4). Then there exists a sufficiently small $\varsigma>0$ for which the following statement holds with probability at least $1-\delta-(\gamma\vartheta)^{\varsigma}$. The final estimate generated by Sp-GD satisfies
$\sum_{j=1}^{k}\left[{\left\lVert\bm{\alpha}_{j}^{t}-\bm{\alpha}_{j}^{\star}\right\rVert}_{2}^{2}+\log^{2}(c_{j}^{t}/c_{j}^{\star})\right]\leq Ck^{3}\pi_{\min}^{-2}\varsigma^{2}\exp\left(\frac{-2\sqrt{\vartheta}\Delta}{\varsigma}\right).$ (43)
The proof of this theorem is deferred to Appendix F.
Remark 4.3.
A few remarks on the statement of Corollary 4.2.
  • Notice that for $\bm{x}$ to satisfy Assumption 2.1, i.e. $\bm{x}$ is subGaussian, $\bm{u}$ needs to be log subGaussian. For example, if $\bm{x}$ is normal, $\bm{u}$ needs to follow the well-known log-normal distribution.
  • For the sake of clarity, the result of Corollary 4.2 is stated where the distortion is only due to RMD. If we assume an additive sub-Gaussian noise model after the RMD transformation, then we can expect an additional error term (43) scaling as $\mathcal{O}(\sigma_{z}^{2}s\log(d)n^{-1})$ similar to the statement of Theorem 2.3. On the other hand, assuming additive sub-Gaussian noise in the generalized polynomial domain requires further investigation.
In summary, Theorem 4.1 and Corollary 4.2 prove that generalized sparse polynomial regression can be done via sparse max-affine regression via the Real Maslov dequantization. We note that such results are the first of their kind. In other words, we are the first to show the approximation (quantization) error for a non-zero choice of $\varsigma>0$ by Theorem 4.1 and a convergence guarantee by Theorem 4.2. The original work by (Maragos et al., 2021) only showed the connection between max-affine functions and (4) in the special case of the posynomial model. Still, this connection is established only at the limit where $\varsigma\to 0$. In other words, the approximation error for a practical choice of $\varsigma$ was not previously shown.
@@ 5.1 Phase Transitions of Sp-GD
This section presents numerical results of the Sp-GD algorithm that corroborate the theoretical guarantees presented in Section 2.2. In the simulation, we initialized Sp-GD in two steps. First, we estimate the parameter subspace by Algorithm 2. Second, we apply a practical alternative to Algorithm 3 which randomly samples from the estimated subspace and chooses the best one that produces the smallest fit error after $10$ iterations of Sp-GD. We adopt this heuristic instead of Algorithm 3 for the following reasons. The first step of Algorithm 3 creates an $r$-covering of the unit $\ell_{2}$-ball $B^{k+1}$, where its cardinality scales as $\mathcal{O}(r^{-k})$ (Vershynin, 2018, Corollary 4.2.12). Therefore, the exhaustive search in the second step over all elements in the $r$-covering becomes impractical as the parameter $r$ decreases. However, the accuracy of Algorithm 3 crucially depends on $r$. The estimation performance is evaluated via the median of the relative error between the true model coefficients $\bm{\theta}^{\star}\triangleq(\bm{\theta_{j}}^{\star})_{j=1}^{k}$ and the estimated coefficients $\widehat{\bm{\theta}}\triangleq(\widehat{\bm{\theta}}_{j})_{j=1}^{k}$ over $50$ Monte Carlo simulations. The relative error is defined via the optimal permutation of the affine model indices as
@@
Figure 1: Median $\mathtt{err}(\widehat{\bm{\theta}}) fo$r different ($n$,$d$) pairs using 50 Monte Carlo iterations for $k=3$ and $s=25$ with Gaussian (top) and Uniform (bottom) covariate distributions in the noiseless case.
@@
Figure 2: Median $\mathtt{err}(\widehat{\bm{\theta}}) fo$r different ($n$,$s$) pairs using 50 Monte Carlo iterations for $k=3$ and $d=200$ with Gaussian (left) and Uniform (right) covariate distributions. The red curves are fitted with respect to $s\log(d/s)$ at the phase transition boundary for both figures.
[…]
@@ 5.2 Subspace Estimation and Random Search for Initialization
This section delves into the detailed empirical analysis of the initialization method employed in the previous section. First, Fig. 4 demonstrates the gain of Algorithm 2 with SPCA over the analogous spectral method with PCA. The regularization parameter $\lambda>0$ required by Algorithm 2 is set by a parameter sweep over a sampling grid. Alternatively, this parameter can be tuned through cross-validation.
@@
Next, we investigate the empirical performance of the repeated random initialization that replaces the exhaustive discrete search in Algorithm 3 by random sampling followed by Sp-GD $10$ iterations. In particular, we observe the estimation accuracy as the function in the number of random initializations $M$. Fig. 5 compares the estimation error by the repeated random initialization when the parameter subspace is estimated by the PCA-based spectral initialization (Ghosh et al., 2021) (red) and by Algorithm 2 (blue). For both of the two subspace estimation methods, the initialization error monotonically decreases as the number of trials $M$ increases. However, the initialization performance by Algorithm 2 is more accurate since it provides better subspace estimation performance.
@@ 6 Discussion
We consider variable selection for a class of nonlinear regression models given by the maximum of $k$ affine models $\bm{x}\mapsto\max_{j\in[k]}\langle\bm{a}_{j}^{\star},\bm{x}\rangle+b_{j}^{\star}$ for $j=1,\dots,k$ where $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ and $\{b_{j}^{\star}\}_{j=1}^{k}$ denote the ground-truth weight vectors and intercepts. The weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ satisfy the joint $s$-sparse structure as we assume that only $s$ out of $d$ covariates contribute to explaining the response variable. This paper proposes a variant of the projected gradient algorithm, Sp-GD, to estimate the sparse model parameters. We provide non-asymptotic local convergence guarantees for Sp-GD under independent sub-Gaussian noise when the covariates follow a sub-Gaussian distribution satisfying the anti-concentration property. Under these assumptions, when the ground-truth model order and parameters are fixed, a suitably initialized Sp-GD converges linearly to an $\epsilon$-accurate parameter estimate given $\mathcal{O}(\max(\epsilon^{-2}\sigma_{z}^{2},1)s\log(d/s))$. In particular, when the observations are noise-free ($\sigma_{z}^{2}=0$), Sp-GD guarantees exact parameter recovery. Since minimizing the squared loss of sparse max-affine models is non-convex, starting Sp-GD within the basin of attraction is crucial for its convergence to the desired estimate. For this purpose, we propose a modification of the spectral method by (Ghosh et al., 2021) which estimates the span of the max-affine weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ so that the subspace estimation utilizes the jointly sparse structure in the weight vectors via sparse principal component analysis. Combined with the $r$-covering search over the estimated subspace, the initialization scheme provides an $\epsilon$-accurate estimate when $r=\mathcal{O}(\epsilon)$ given $\mathcal{O}(\epsilon^{-2}\max(\sigma_{z}^{4},\sigma_{z}^{2},1)s^{2}\log^{4}d)$ observations when the ground-truth model parameters are fixed, and the covariates and noise follow Gaussian distributions. The dominating factor of the sample complexity is $s^{2}$, which is significantly smaller than $d$ in the non-sparse case.
@@
One noteworthy limitation of the initialization is the $r$-covering search inherited from the previous work on non-sparse max-affine regression (Ghosh et al., 2021). The cost of constructing the $r$-covering and the exhaustive search over it increases exponentially in the subspace dimension. Therefore, the $r$-covering search is not practical for small separation parameter $r$ which is necessary for accurate estimates. However, this is the only known algorithm with theoretical guarantees for max-affine parameter initialization. Therefore, it would be a fruitful future direction to develop a practical initialization scheme that avoids the $r$-covering search and provides theoretical performance guarantees. Alternatively, it would also be intriguing to explore a potential analysis of Sp-GD from random initialization that will extend the known theoretical results on single-index models (Tan and Vershynin, 2019; Chandrasekher et al., 2022).
@@
$\left\|\bm{\theta}^{t+1}-\bm{\theta}^{\star}\right\|^{2}_{2}\leq\tau\left\|\bm{\theta}^{t}-\bm{\theta}^{\star}\right\|^{2}_{2}+C_{1}\sigma^{2}_{z}\left(\frac{sk\log\left(n/s\right)+s\log\left({d}/{s}\right)+\log\left({1}/{\delta}\right)}{n}\right),$ (44a)
$\bm{\theta}^{t+1}\in\mathcal{N}(\bm{\theta}^{\star},\sqrt{2}\Delta\rho),$ (44b)
hold for all $t\in\mathbb{N}\cup\{0\}$. Here $\delta$ refers to the error probability, i.e. these statements hold with probability at least $1-\delta$. We prove this statement by induction. Let $t$ be arbitrarily fixed. We assume that (44) holds for all previous iterates. To show that (44a) also holds for the current iterate, we introduce the following notation. Let $\mathcal{S}^{\star}$ and $\mathcal{S}^{t}$ denote the joint support of $\{\bm{\theta}_{j}^{\star}\}_{j=1}^{k}$ and $\{\bm{\theta}_{j}^{t}\}_{j=1}^{k}$, respectively. The union of $\mathcal{S}^{\star}$ and $\mathcal{S}^{t}$ is denoted by $\mathcal{U}^{t}$. We use $\bm{\Pi}_{\mathcal{U}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to denote the orthogonal projection onto the subspace spanned by the standard basis vectors $\{\bm{e}_{j}\}_{j\in\mathcal{U}}$ for $\mathcal{U}\subset[d]$. Then the augmented operator $\bm{\widetilde{\Pi}}_{\mathcal{U}}:\mathbb{R}^{d+1}\rightarrow\mathbb{R}^{d+1}$ is defined as
$\bm{\widetilde{\Pi}}_{\mathcal{U}}=\left[{\begin{array}[]{cc}\bm{\Pi}_{\mathcal{U}}&\bm{0}_{d\times 1}\\ \bm{0}_{1\times d}&1\\ \end{array}}\right].$ (45)
The step size used in Sp-GD stated in Algorithm 1 is stated as
@@
$\left\|\bm{\theta}^{t+1}-\bm{\theta}^{\star}\right\|_{2}$ $=\left\|\left(\bm{I}_{k}\otimes\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\right)\left(\bm{\theta}^{t+1}-\bm{\theta}^{\star}\right)\right\|_{2}$
$\leq\left\|\left(\bm{I}_{k}\otimes\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\right)\left(\bm{\theta}^{t+1}-\bm{\alpha}^{t+1}\right)\right\|_{2}+\left\|\left(\bm{I}_{k}\otimes\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\right)\left(\bm{\alpha}^{t+1}-\bm{\theta}^{\star}\right)\right\|_{2}$
$=\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{\theta}^{t+1}_{j}-\bm{\alpha}^{t+1}_{j}\right)\right\|_{2}^{2}\right)^{1/2}+\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{\alpha}^{t+1}_{j}-\bm{\theta}^{\star}_{j}\right)\right\|_{2}^{2}\right)^{1/2}$
$\leq 2\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{\alpha}^{t+1}_{j}-\bm{\theta}^{\star}_{j}\right)\right\|_{2}^{2}\right)^{1/2}$
$=2\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{\theta}^{t}_{j}-\mu_{j}^{t}\nabla_{\bm{\theta}_{j}}\ell(\bm{\theta}^{t})-\bm{\theta}^{\star}_{j}\right)\right\|_{2}^{2}\right)^{1/2},$ (47)